OpenStudy (anonymous):

Write the first five terms of the geometric sequence. an = -7an-1; a1 = -2

OpenStudy (anonymous):

\[a_n = -7*a_{n-1}\] replace 2 into n and you'll get that \[a_2 = -7*a_{1} = -7*-2 = 14\]

OpenStudy (anonymous):

-2 14 ?? ?? ?? In case you haven't realized the ratio you're constantly multiplying by is -7